OpenStudy (anonymous):

What is the equation of the line, in general form, that passes through point (-1, -1) and is parallel to the line whose equation is x + y = 3? a. x - y + 2 = 0 b. x+ y - 2 = 0 c. x + y + 2 = 0

OpenStudy (anonymous):

can you tell me what the slope is to the line given?

OpenStudy (anonymous):

parallel lines have same slope... y - y1 = m(x -x1) x1,y1 given point m = you are calculating